# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v1

@@ 5.1 Numerical results
5.1.1 Experiment with synthetic data
@@
To observe the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, and ${k}_{n}^{\operatorname{sep},q}$ with different values of $n$ and compare them to the existing kernels $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ (corresponding to $n=\infty$), we consider a regression task with synthetic data. For $i=1,\ldots,N$, we generated input training samples as $x^{i}(z)=[\sin(0.01iz)+0.01\xi_{z,i},\cos(0.01iz)+0.01\eta_{z,i}]$. We set the target function $f$ as $f(x)(z)=3(\sin(\cos(\int_{z-\Delta}^{z+\Delta}x_{1}(t)\mathrm{d}t+\int_{z-\Delta}^{z+\Delta}x_{2}(t)\mathrm{d}t)))$, where $\Delta=2\pi/30$, and generated training output samples $y^{i}(z)=f(x)(z)+0.001\xi_{z,i}$. Here, $\xi_{z,i}$ and $\eta_{z,i}$ are sampled independently from the Gaussian distribution with mean 0 and standard deviation 1. We set the sample size $N$ as 1000. For the kernel ${k}_{n}^{\operatorname{poly},q}$, we set $q=1$ and $\alpha_{1}=1$. Note that ${k}^{\operatorname{poly},q}$ is the linear kernel in this case. For the kernel $\hat{k}_{n}^{\operatorname{prod},q}$, we set $q=1$, $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=\mathrm{e}^{-|x-y|^{2}}$. We set $\beta_{n}=1$ for $n<\infty$ and $\beta_{\infty}=0$. For the kernel $k_{n}^{\operatorname{sep},q}$, we set $\tilde{k}(x,y)=\mathrm{e}^{-0.1/\Delta^{2}(\|x_{1}-y_{1}\|_{L^{2}(\mathbb{T})}^{2}+\|x_{2}-y_{2}\|_{L^{2}(\mathbb{T})}^{2})}$, $q=2$, and $a_{1}(z)=a_{2}(z)=\mathrm{e}^{\sin z}$. We estimated $f$ using kernel ridge regression. Figure 2 (a-c) illustrates the test error. We can see that for ${k}_{n}^{\operatorname{poly},q}$, the test error is the smallest when $n=32$. It becomes large if $n$ is larger or smaller than $32$. This is because when $n=32$, the empirical error and the complexity in Eq. 3 are balanced, and the best possible expected error is obtained. We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the existing kernel ($n=\infty$). We have similar results for the other two kernels. Note that $\hat{k}_{\infty}^{\operatorname{prod},q}(x,y)(z)=\hat{k}^{\operatorname{prod},q}(x,y)(z)$ only depends on $x(z)$ and $y(z)$. However, $f(x)(z)$ depends on $x(t)$ for $t\in[z-\Delta,z+\Delta]$. Thus, the test error becomes large when $n=\infty$. In addition, as discussed in Remark 14, $\hat{k}^{\operatorname{prod},q}$ may not be positive definite. However, although we set $\beta_{n}$ as a small value, all the eigenvalues of the Gram matrix are positive in this case. See Appendix D.1 for more details.
@@
Figure 2: (a-c) Test error of the regression task with different values of $n$. (Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (d) Test error of the image recovering task with different values of $n$. (Box plot of results of five independent runs with different training and test data.) (e) Output images of the image recovering task with different values of $n$.
5.1.2 Experiment with MNIST
We consider an image recovering task with MNIST [30]. For $i=1,\ldots,N$, we generated input training samples by setting all the $8\times 8$ pixels in the middle of the images as 00. We set the output samples as the original images. We flatten the 2-dimensional $28\times 28$ image to a vector in $\mathbb{R}^{784}$ and regard it as a discretized function on the space of pixels. We try to estimate a function that transforms an image with the missing part into the original image. We applied kernel ridge regression with the regularization parameter $\lambda=0.01$. We set the sample size $N$ as $200$, and used the kernel $k_{n}^{\operatorname{prod},q}$ with $q=1$ and $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=\mathrm{e}^{-0.1|x-y|^{2}}$. We set $\beta_{n}=0.01$ for $n<\infty$ and $\beta_{\infty}=0$. Figure 2 (d) and (e) show the test error and the output images with different values of $n$. We can see that the test error becomes the smallest when $n=70$, but it becomes large when $n$ is smaller or larger than $70$. As for the output images, when $n=\infty$, i.e., the existing kernel $k^{\operatorname{prod},q}$, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image. When $n=70$, we can recover the image the clearest.
6 Generalization to Other $C^{*}$-algebras
Setting approximate order isomorphisms for more general $C^{*}$-algebras has been investigated. Using these results and replacing $R_{n}$ and $S_{n}$ in Definition 6, we can define positive definite kernels for the $\mathcal{A}^{d}$-valued inputs and $\mathcal{A}$-valued output for a more general $C^{*}$-algebra $\mathcal{A}$.
Continuous functions on high-dimensional torus [20]
[…]
Continuous functions on the sphere [19]
@@ 7 Conclusion and Limitation
In this paper, we proposed a new class of positive definite kernels for functional inputs and outputs based on the spectral truncation. The proposed kernels are indexed by a truncation parameter $n$, and they converge to existing kernels: polynomial, product, and separable kernels, as $n$ goes to infinity. Setting $n$ as a finite value, we can introduce noncommutativity into the learning process. More precisely, we can induce interactions along the data function domain and control the tradeoff between the representation power and the model complexity of the RKHMs associated with the proposed kernels. Since $n$ describes the resolution of the discretization functions, our results show the positive effect of discretization on generalization. We applied the proposed kernel to kernel ridge regression and numerically observed that the proposed kernels outperformed the existing kernels.
In the current setting, we focus on the $C^{*}$-algebra of continuous functions on the torus. Although we can generalize this setting and define positive definite kernels for more general inputs and outputs as we discussed in Section 6, theoretical investigations for these generalized settings and their computationally efficient implementations remain to be investigated in future work.
@@ Appendix A Proofs
Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)\to k^{\operatorname{poly},q}(x,y)$. The proofs for $k_{n}^{\operatorname{prod},q}(x,y)$ and $k_{n}^{\operatorname{sep},q}(x,y)$ are the same. We have
$k_{n}^{\operatorname{poly},q}(x,y)=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}$
@@
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j=0}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{q}}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
@@
where the sum $\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{-1}$ is set as 00, $g_{i}(t)=\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$ for $t=[t_{1},\ldots,t_{2q}]\in\mathbb{T}^{2q}$, and $P=\{r=[r_{1},\ldots,r_{2q}]\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{m}r_{i}|\leq 1,\ l\leq m\}$. In addition, we set $j=r_{0}$ and $j=r_{q}$. The second to the last equality is derived by Lemma 8. Since $P$ is a convex polyhedron, the last equality is by Lemma 9. For $k_{n}^{\operatorname{sep},q}$, we additionally use the fact $\lim_{n\to\infty}\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))=\tilde{k}(\lim_{n\to\infty}S_{n}(R_{n}(x)),\lim_{n\to\infty}S_{n}(R_{n}(y)))$, which follows from the continuity of $\tilde{k}$, and the fact $S_{n}(R_{n}(x))\to x$ uniformly as $n\to\infty$ if $x$ and $y$ are differentiable [17, Lemma 10]. $\Box$
@@
Proof  Let $r^{\prime}\in Q_{j}^{m}$. For $i>j$, we have $r_{i}=r_{i}^{\prime}+r_{i-1}=r_{i}^{\prime}+(r_{i-1}^{\prime}+r_{i-2})=\cdots=m+r_{j+1}^{\prime}+\cdots+r_{i}^{\prime}$, and for $i<j$, we have $-r_{i}=r_{i+1}^{\prime}-r_{i+1}=r_{i+1}^{\prime}+(r_{i+2}^{\prime}-r_{i+2})=\cdots=r_{i+1}^{\prime}+\cdots+r_{j}^{\prime}-m$. Therefore, we have
@@
Proposition 11    The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep}}$ are positive definite.
@@
Proposition 12    Let $\beta_{n}\geq-\min_{t\in\mathbb{T}^{q}}F_{n}^{q,P}(t)$. Let
@@
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}k_{n}^{\operatorname{prod},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)F_{n}^{2q,P}(z\mathbf{1}-t)\mathrm{d}t$
$\qquad+\beta_{n}\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)\mathrm{d}t$
$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\prod_{j=1}^{q}\overline{d_{m}(z)}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))d_{l}(z)(F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n})\mathrm{d}t$
@@
for $z\in\mathbb{T}$. The last inequality is derived since the map $([x_{1},x_{2}],[y_{1},y_{2}])\mapsto\overline{\tilde{k}_{1,j}(x_{1},y_{1})}\tilde{k}_{2,j}(x_{2},y_{2})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$. $\Box$
@@
Proposition 16    Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class. Assume for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have $\hat{R}_{N}(\mathbf{x},\mathcal{F}(z))=\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})(z)$, where $\mathcal{F}(z)=\{x\mapsto(f(x))(z)\,\mid\,f\in\mathcal{F}\}$.
@@
Theorem 15    Assume $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep}q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
According to Hashimoto et al. [14], the $\mathcal{A}$-valued Rademacher complexity for the RKHM associated with an $\mathcal{A}$-valued kernel $k$ is upperbounded by $\|k(x_{i},x_{i})\|_{\mathcal{A}}$ as follows.
@@
Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq C\}$. Then, we have
We obtain the following lemma directly by Lemma 4.2 in [31].
Lemma 18
Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. Assume also for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have
@@
Let $g$ be the same map defined in Lemma 18. For any $f\in\mathcal{F}$ and any $\delta\in(0,1)$, with probability at least $1-\delta$, we have

### v3

@@ 7 Numerical results
7.1 Experiment with synthetic data
@@
We first observed the behavior of the proposed kernels $k_{n}^{\operatorname{poly},q}$, $\hat{k}_{n}^{\operatorname{prod},q}$, and ${k}_{n}^{\operatorname{sep},q}$ with different values of $n$ and compared them to the commutative kernels $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ (corresponding to $n=\infty$). We considered a regression task with synthetic data. For $i=1,\ldots,N$, we generated input training samples as $x^{i}(z)=[\sin(0.01iz)+0.01\xi_{z,i},\cos(0.01iz)+0.01\eta_{z,i}]$. We set the target function $f$ as $f(x)(z)=3(\sin(\cos(\int_{z-\Delta}^{z+\Delta}x_{1}(t)\mathrm{d}t+\int_{z-\Delta}^{z+\Delta}x_{2}(t)\mathrm{d}t)))$, where $\Delta=2\pi/30$, and generated training output samples $y^{i}(z)=f(x)(z)+0.001\xi_{z,i}$. Here, $\xi_{z,i}$ and $\eta_{z,i}$ were sampled independently from the Gaussian distribution with mean 0 and standard deviation 1. We set the sample size $N$ as 1000. For the kernel ${k}_{n}^{\operatorname{poly},q}$, we set $q=1$ and $\alpha_{1}=1$. Note that ${k}^{\operatorname{poly},q}$ is the linear kernel in this case. For the kernel $\hat{k}_{n}^{\operatorname{prod},q}$, we set $q=1$, $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=2\pi\mathrm{e}^{-|x-y|^{2}}$. We set $\beta_{n}=1$ for $n<\infty$ and $\beta_{\infty}=0$. For the kernel $k_{n}^{\operatorname{sep},q}$, we set $\tilde{k}(x,y)=\mathrm{e}^{-0.1\cdot 2\pi^{2}(\|x_{1}-y_{1}\|_{L^{2}(\mathbb{T})}^{2}+\|x_{2}-y_{2}\|_{L^{2}(\mathbb{T})}^{2})}$, $q=2$, and $a_{1}(z)=a_{2}(z)=2\pi\mathrm{e}^{\sin z}$. We estimated $f$ using kernel ridge regression. We applied the cross-validation grid search to find an optimal regularization parameter $\lambda$. Figure 2 (a-c) illustrates the test error. We can see that for ${k}_{n}^{\operatorname{poly},q}$, the test error is the smallest when $n=32$. It becomes large if $n$ is larger or smaller than $32$. This is because when $n=32$, the empirical error and the complexity in Eq. 3 are balanced, and the best possible expected error is obtained. We can see that even for the simplest kernel ${k}_{n}^{\operatorname{poly},1}$, the proposed kernel ($n$ is finite) goes beyond the typical commutative kernel ($n=\infty$). We have similar results for the other two kernels. Note that $\hat{k}_{\infty}^{\operatorname{prod},q}(x,y)(z)=\hat{k}^{\operatorname{prod},q}(x,y)(z)$ only depends on $x(z)$ and $y(z)$. However, $f(x)(z)$ depends on $x(t)$ for $t\in[z-\Delta,z+\Delta]$. Thus, the test error becomes large when $n=\infty$. In addition, as discussed in Remark 15, $\hat{k}_{n}^{\operatorname{prod},q}$ may not be positive definite. However, although we set $\beta_{n}$ as a small value, all the eigenvalues of the Gram matrix are positive in this case. See Appendix E.1 for more details.
We also compared the performance and the computational time of RKHMs associated with the proposed kernels to a vvRKHS with the same setting as above. For the vvRKHS, we used the nonseparable kernel $k(x,y)=\tilde{k}(x,y)[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$, where we replaced $S_{n}(\prod_{j=1}^{q}R_{n}(a_{j})^{*}\prod_{j=1}^{q}R_{n}(a_{j}))$ with the nonseparable matrix $[\mathrm{e}^{-|x_{i}-y_{j}|^{2}}]_{i,j}$. This nonseparable kernel is proposed by Lim et al. (2015). Table 1 shows the result. We can see that the RKHMs outperform the vvRKHS, and the computational time for the RKHMs are smaller than that for the vvRKHS. Note that we have already confirmed the proposed kernels outperform the existing separable kernel $k_{\infty}^{\operatorname{sep},q}$ in the above experiment (See Figure 2 (c)).
Finally, we observed the behavior of the deep model. We generated data in the same manner as above, but we set $N=300$. We used the kernel ${k}_{n}^{\operatorname{prod},q}$ and set $L=1,2,3$, $\tau_{1}=\sqrt{2}$, $\tau_{2}=\sqrt{3}$, and $\tau_{3}=\sqrt{5}/2$. We set $n$ so that the numbers of parameters are the same for all values of $L$. We used the loss function $\|\sum_{i=1}^{N}|f(x_{i})-y_{i}|_{k}\|_{L^{2}(\mathbb{T})}+0.1\||f|_{k}\|_{L^{2}(\mathbb{T})}$ for the deep model $f$ defined in Eq. (6). Figure 2 (d) shows the result. We can see that as the number of layers increases, we can obtain a higher performance model even with the same number of parameters.
@@
Figure 2: (a-c) Test error of the regression task with different values of $n$. (Box plot of results of five independent runs with different values of noises $\xi_{z,i}$ and $\eta_{z,i}$ in $x^{i}(z)$ and $y^{i}(z)$.) (d) Test error of the regression task with deep approach with different $n$ and $L$. The parameters $n$ and $L$ are chosen so that the numbers of parameters are the same for all the cases. (e) Output images of the image recovering task with different values of $n$.
Table 1: Test error and CPU time of the regression task with RKHM and vvRKHS
Test Error CPU Time (s)
RKHM, $k=k_{n}^{\operatorname{prod},q}$ ($n=9$) 0.0113$\pm$1.63×$\times 10^{-5}$ 149.3$\pm$2.392
RKHM, $k=k_{n}^{\operatorname{sep},q}$ ($n=10$) 0.0385$\pm$5.04388×$\times 10^{-5}$ 22.79$\pm$ 0.5742
vvRKHS 0.0774$\pm$7.86 $10^{-5}$ 570.4$\pm$14.87
7.2 Experiment with MNIST
We consider an image recovering task with MNIST (LeCun et al., 1998). For $i=1,\ldots,N$, we generated input training samples by setting all the $8\times 8$ pixels in the middle of the images as 00. We set the output samples as the original images. We flatten the 2-dimensional $28\times 28$ image to a vector in $\mathbb{R}^{784}$ and regard it as a discretized function on the space of pixels. We tried to estimate a function that transforms an image with the missing part into the original image. We applied kernel ridge regression with the regularization parameter $\lambda=0.01$. We set the sample size $N$ as $200$, and used the kernel $k_{n}^{\operatorname{prod},q}$ with $q=1$ and $\tilde{k}_{1,1}(x,y)=\tilde{k}_{2,1}(x,y)=2\pi\mathrm{e}^{-0.1|x-y|^{2}}$. We set $\beta_{n}=0.01$ for $n<\infty$ and $\beta_{\infty}=0$. Figure 2 (e) shows the output images with different values of $n$. When $n=\infty$, i.e., the commutative kernel, we cannot recover the missing part since each pixel of the output is determined only with the corresponding pixel of the input, and we cannot obtain information of other pixels to estimate what is written in the image. When $n=70$, we can recover the image the clearest.
@@ 8 Conclusion and Limitation
In this paper, we proposed a new class of positive definite kernels for vector- or function-valued outputs based on the spectral truncation. The proposed kernels with the framework of RKHMs resolve two shortcomings of the framework of vvRKHSs for vector- or function-valued outputs at the same time: computational cost and choice of kernels. By considering function-valued positive definite kernels, not operator-valued ones, we can alleviate the computational cost. At the same time, we can introduce noncommutativity into the learning process and can induce interactions along the data function domain. We also showed that we can control the tradeoff between the representation power and the model complexity associated with the proposed kernels. In addition, we proposed a deep approach to obtain models with higher representation powers.
In the current setting, we focus on the $C^{*}$-algebra of continuous functions on the torus. Although this setting includes many important examples, e.g., periodic time-series data, continuous functions defined on a compact set in $\mathbb{R}$, finite-dimensional vectors, investigations for more general $C^{*}$-algebras allow us to apply the proposed kernel more general settings. As we discuss in Appendix D, we can generalize this setting and define positive definite kernels for more general inputs and outputs. However, theoretical investigations for these generalized settings and their computationally efficient implementations remain to be investigated in future work.
@@ Appendix A Proofs
Proof  We show $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$. The proofs for $k_{n}^{\operatorname{prod},q}(x,y)$ and $k_{n}^{\operatorname{sep},q}(x,y)$ are similar. We have
$k_{n}^{\operatorname{poly},q}(x,y)(z)=S_{n}\bigg{(}\sum_{i=1}^{d}\alpha_{i}(R_{n}(x_{i})^{*})^{q}R_{n}(y_{i})^{q}\bigg{)}(z)$
@@
$=\frac{1}{n}\sum_{i=1}^{d}\alpha_{i}\sum_{j=0}^{n-1}\sum_{r\in jP\bigcap\mathbb{Z}^{2q}}\int_{\mathbb{T}}\cdots\int_{\mathbb{T}}\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$
@@
where the sum $\sum_{j,l,r_{1},\ldots,r_{2q-1}=0}^{-1}$ is set as 00, $g_{i}(t)=\overline{x_{i}(t_{1})}\cdots\overline{x_{i}(t_{q})}y_{i}(t_{q+1})\cdots y_{i}(t_{2q})$ for $t=[t_{1},\ldots,t_{2q}]\in\mathbb{T}^{2q}$, and $P=\{r=[r_{1},\ldots,r_{2q}]\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{m}r_{i}|\leq 1,\ l\leq m\}$. In addition, we set $j=r_{0}$ and $j=r_{q}$. The second to the last equality is derived by Lemma 9 below. Since $P$ is a convex polyhedron, $g_{i}\ast F_{n}^{2q,P}(z)\to g_{i}(z)$ as $n\to\infty$ by Lemma 10. For $k_{n}^{\operatorname{sep},q}$, we additionally use the fact $\lim_{n\to\infty}\tilde{k}(S_{n}(R_{n}(x)),S_{n}(R_{n}(y)))=\tilde{k}(\lim_{n\to\infty}S_{n}(R_{n}(x)),\lim_{n\to\infty}S_{n}(R_{n}(y)))$, which follows from the continuity of $\tilde{k}$, and the fact $S_{n}(R_{n}(x))\to x$ uniformly as $n\to\infty$ if $x$ and $y$ are differentiable (van Suijlekom, 2021, Lemma 10). $\Box$
@@
Proof  Let $r^{\prime}\in Q_{j}^{m}$. For $i>j$, we have $r_{i}=r_{i}^{\prime}+r_{i-1}=r_{i}^{\prime}+(r_{i-1}^{\prime}+r_{i-2})=\cdots=r_{i}^{\prime}+\cdots+r_{j+1}^{\prime}+m$, and for $i<j$, we have $-r_{i}=r_{i+1}^{\prime}-r_{i+1}=r_{i+1}^{\prime}+(r_{i+2}^{\prime}-r_{i+2})=\cdots=r_{i+1}^{\prime}+\cdots+r_{j}^{\prime}-m$. Therefore, we have
@@
Proposition 12    The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite.
@@
Proposition 13    Let $\beta_{n}\geq-\min_{z\in\mathbb{T}^{q}}F_{n}^{2q,P}(z)$. Then, $\hat{k}_{n}^{\operatorname{prod},q}$ defined below is positive definite.
@@
$\bigg{(}\sum_{j,l=1}^{N}d_{j}^{*}\hat{k}_{n}^{\operatorname{prod},q}(x_{j},x_{l})d_{l}\bigg{)}(z)$
$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\overline{d_{m}(z)}\bigg{(}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))\bigg{)}d_{l}(z)F_{n}^{2q,P}(z\mathbf{1}-t)\mathrm{d}t$
$\qquad+\beta_{n}\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\overline{d_{m}(z)}\bigg{(}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))\bigg{)}d_{l}(z)\mathrm{d}t$
$=\sum_{m,l=1}^{N}\int_{\mathbb{T}^{2q}}\overline{d_{m}(z)}\bigg{(}\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{m}(t_{j}),x_{l}(t_{j}))}\tilde{k}_{2,j}(x_{m}(t_{q+j}),x_{l}(t_{q+j}))\bigg{)}d_{l}(z)(F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n})\mathrm{d}t$
@@
for $z\in\mathbb{T}$. The last inequality is derived since the map $([x_{1},\ldots,x_{2q}],[y_{1},\ldots,y_{2q}])\mapsto\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x_{2j-1},y_{2j-1})}\tilde{k}_{2,j}(x_{2j},y_{2j})$ is positive definite and $F_{n}^{2q,P}(z\mathbf{1}-t)+\beta_{n}\geq 0$. $\Box$
@@
Let $\Omega$ be a probability space with a probability measure $\mu$. Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be samples from a distributions of $\mathcal{A}_{0}^{d}$-valued random variable $X$ and $\mathcal{A}_{1}$-valued random variable $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$). Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$.
Theorem 16    Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep}q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|_{\operatorname{op}}^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
To show Theorem 16, we use the $\mathcal{A}$-valued Rademacher complexity defined by Hashimoto et al. (2023a). Let $\sigma_{1},\ldots,\sigma_{N}$ be i.i.d. Rademacher variables, which take their values on $\{-1,1\}$. For a real-valued function class $\mathcal{F}$ and $\mathbf{x}=[x_{1},\ldots,x_{N}]\in(\mathcal{A}^{d})^{N}$, the empirical Rademacher complexity $\hat{R}_{N}(\mathbf{x},\mathcal{F})$ is defined by $\hat{R}_{N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}f(x_{i})\sigma_{i}]/N$. Here, $\mathrm{E}$ is the integration on $\Omega$ with respect to $\mu$. Similar to the case of the standard Rademacher complexity, an $\mathcal{A}$-valued version of the Rademacher complexity is defined as $\hat{R}_{\mathcal{A},N}(\mathbf{x},\mathcal{F})=\mathrm{E}[\sup^{\mathcal{A}}_{f\in\mathcal{F}}\sum_{i=1}^{N}|f(x_{i})^{*}\sigma_{i}|_{\mathcal{A}}]/N$ for an $\mathcal{A}$-valued function class $\mathcal{F}$. Here, $|a|_{\mathcal{A}}=(a^{*}a)^{1/2}$ for $a\in\mathcal{A}$, $\sup^{\mathcal{A}}$ is the supremum in the sense of the order in $\mathcal{A}$ (see Definition 3), and the integral $\mathrm{E}$ means the Bochner integral in this case.
Indeed, in our case, the $\mathcal{A}$-valued Rademacher complexity is represented by the standard Rademacher complexity, and the argument is reduced to evaluate the standard Rademacher complexity. In the following, we denote by $\mathcal{A}_{+}$ the subset of $\mathcal{A}$ composed of all positive elements in $\mathcal{A}$.
Proposition 20
@@
We obtain the following lemma directly by Lemma 4.2 in Mohri et al. (2012).
@@
Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class. Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. Assume also for any $x\in\mathcal{A}_{0}^{d}$ and $z\in\mathbb{T}$, $f(x)(z)\in\mathbb{R}$. Then, we have
@@
Let $\mathcal{F}$ be an $\mathcal{A}$-valued function class. Let $g$ be the same map defined in Lemma 21. For any $f\in\mathcal{F}$ and any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
Here, $\mathrm{E}_{\sigma_{i}}[\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}]=\int_{\Omega}\sup_{f\in\mathcal{F}}\sum_{i=1}^{N}g(f(X_{i})(z),Y_{i}(z))\sigma_{i}(\omega)\mathrm{d}\mu(\omega)$.